OpenStudy (anonymous):

The sequence is defined by defining the a_i to be ai=(-1)^i+1(1/3)^i What is a_2 ?

OpenStudy (tkhunny):

How did you get into a class with this kind of material and manage to miss everything about the Order of Operations? Do you mean this? \(a_{i} = (-1)^{i+1}(1/3)^{i}\) If so, why not give writing it correctly another go. You may need a lesson on that. Use parentheses to clarify meaning.

OpenStudy (anonymous):

I didn't choose to be in Caucus:( they put me in here making me think it's algebra 2.. I did mean the equation you wrote, I am just unable to type it that way through my computer.

OpenStudy (anonymous):

omg.. my spelling... calculus.. sorry.

OpenStudy (anonymous):

ai=(−1)i+1(1/3)i Does this work?

OpenStudy (tkhunny):

You didn't understand what I said. You have written \(a_{i} = (-1)\cdot i + 1\cdot (1/3)\cdot i\). I seriously doubt this is what you intend. The karat usually symbolizes exponentiation. \((-i)^{i+1}\) = (-1)^(i+1) For example. Use parentheses to clarify intent. Write so that what you mean is the most likely interpretation.

OpenStudy (anonymous):

I am just going to skip this question on my study guide.. Thank anyways!

OpenStudy (tkhunny):

Why not learn to communicate? If you mean \(a_{i} = (-1)^{i+1}\cdot (1/3)^{i}\), just write that. Perhaps this: ai = (-1)^(i+1) * (1/3)^i Now, substitute i = 2.

OpenStudy (anonymous):

I have a learning disability in math, so it's a little harder for me understand and figure out how to communicate when it comes to math.. I am trying, I have substituted the 2 and Im just lost.

OpenStudy (tkhunny):

2+1 = 3 \(a_{2} = (-1)^{3}\cdot (1/3)^{2} = (-1)\cdot (1/9) = -\dfrac{1}{9}\) Are you SURE there is something in there that you can't do?